I neeed help with the following problems :
1.) 51/d = 39/14
2.) 29/77 = b/78
3.) 62/2 = c/57
4.) 67/b = 17/42
5.) 61/b = 41/49

Answer by checkley79(3341)

1.) 51/d = 39/14
39d=51*14
39d=714
d=714/39
d=18.31 ans
2.) 29/77 = b/78
77b=29*78
77b=2262
b=2262/77
b=29.38 ans.
3.) 62/2 = c/57
2c=62*57
2c=3534/2
c=1767 ans.
4.) 67/b = 17/42
17b=67*42
17b=2814
b=1814/17
b=165.53 ans.
5.) 61/b = 41/49
41b=61*49
41b=2989
b=2989/41
b=72.9 ans.